# DualOptim: Enhancing Efficacy and Stability in Machine Unlearning with Dual Optimizers

### v1

@@ B.1 Implementation Details for Image Classification
For CIFAR-10, CIFAR-100, and SVHN using ResNet-18, all baselines use the SGD optimizer with momentum of $0.9$, weight decay of $5\times 10^{-4}$, and batch size of $128$ if not specified. For TinyImageNet, Swin-T, the models are initialized from torchvision weight pre-trained on ImageNet. All baselines use the Adam optimizer with a weight decay of $5\times 10^{-4}$ and batch size of $128$ if not specified. Summaries of the hyperparameters for each method on each dataset are shown in Table 6 - 9. Note that the unspecified hyperparameters are the same as the default ones reported in their original papers.
@@
Pretrain epoch $=200$, cosine scheduler, $\eta=0.1$
RT epoch $=200$, cosine scheduler, $\eta=0.1$
FT epoch $=10$, constant scheduler, $\eta=0.01$
GA epoch $=10$, constant scheduler, $\eta=5\times 10^{-4}$
RL epoch $=10$, constant scheduler, $\eta_{f}=0.01$, $\eta_{r}=0.01$
SCRUB epoch $=10$, constant scheduler, Adam, $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch $=10$, constant scheduler, $\eta_{f}=0.018$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch $=10$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=1.5\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch $=10$, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=1.3\times 10^{-4}$, $\eta_{r}=0.01$
@@
Pretrain epoch $=200$, cosine scheduler, $\eta=0.1$
RT epoch $=200$, cosine scheduler, $\eta=0.1$
FT epoch $=10$, constant scheduler, $\eta=0.01$
GA epoch $=10$, constant scheduler, $\eta=8\times 10^{-5}$
RL epoch $=10$, constant scheduler, $\eta_{f}=0.01$, $\eta_{r}=0.01$
SCRUB epoch $=10$, constant scheduler, Adam, $\eta_{f}=1.6\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch $=10$, constant scheduler, $\eta_{f}=2.5\times 10^{-4}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch $=10$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2.8\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch $=10$, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=5.5\times 10^{-5}$, $\eta_{r}=0.01$
@@
Pretrain epoch $=200$, cosine scheduler, $\eta=0.1$
RT epoch $=200$, cosine scheduler, $\eta=0.1$
FT epoch $=10$, constant scheduler, $\eta=4.5\times 10^{-2}$
GA epoch $=10$, constant scheduler, $\eta=5.2\times 10^{-4}$
RL epoch $=10$, constant scheduler, $\eta_{f}=8\times 10^{-4}$, $\eta_{r}=0.01$
SCRUB epoch $=10$, constant scheduler, Adam, $\eta_{f}=6\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch $=10$, constant scheduler, $\eta_{f}=1.3\times 10^{-3}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch $=10$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch $=10$, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=1.9\times 10^{-4}$, $\eta_{r}=0.01$
@@
Pretrain epoch $=200$, cosine scheduler, $\eta=0.1$
RT epoch $=200$, cosine scheduler, $\eta=0.1$
FT epoch $=10$, constant scheduler, $\eta=2.2\times 10^{-2}$
GA epoch $=10$, constant scheduler, $\eta=5.2\times 10^{-4}$
RL epoch $=10$, constant scheduler, $\eta_{f}=8\times 10^{-3}$, $\eta_{r}=0.01$
SCRUB epoch $=10$, constant scheduler, Adam, $\eta_{f}=8\times 10^{-6}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch $=10$, constant scheduler, $\eta_{f}=1.25\times 10^{-2}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch $=10$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2.2\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch $=10$, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=2.5\times 10^{-6}$, $\eta_{r}=0.01$
@@
Pretrain epoch $=10$, cosine scheduler, $\eta=1\times 10^{-4}$
RT epoch $=10$, cosine scheduler, $\eta=1\times 10^{-4}$
FT epoch $=1$, constant scheduler, $\eta=1\times 10^{-4}$
GA epoch $=1$, constant scheduler, $\eta=5\times 10^{-6}$
RL epoch $=1$, constant scheduler, $\eta_{f}=5\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SCRUB epoch $=1$, constant scheduler, Adam, $\eta_{f}=2\times 10^{-5}$, $\eta_{r}=2\times 10^{-5}$
SalUn epoch $=1$, constant scheduler, $\eta_{f}=9\times 10^{-5}$, $\eta_{r}=9\times 10^{-5}$
@@
SCRUB+DO epoch $=1$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2\times 10^{-5}$, $\eta_{r}=2\times 10^{-5}$
SalUn+DO epoch $=1$, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=9\times 10^{-5}$, $\eta_{r}=9\times 10^{-5}$
@@ B.3 Implementation Details for Natural Language Processing
Phi-1.5 ME+GD+DO epoch $=5$, $\eta_{f}=1\times 10^{-5}$, $\eta_{r}=1\times 10^{-5}$
IDK+AP+DO epoch $=5$, $\eta_{f}=8\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
LLaMA 2 ME+GD+DO epoch $=5$, $\eta_{f}=1.45\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
IDK+AP+DO epoch $=5$, $\eta_{f}=1.25\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
LLaMA 2 LoRA ME+GD+DO epoch $=5$, $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=2\times 10^{-4}$
IDK+AP+DO epoch $=5$, $\eta_{f}=5\times 10^{-5}$, $\eta_{r}=2\times 10^{-4}$
@@ Appendix C Additional Results
C.1 Low Correlation between Forget and Retain Graidients
@@
Figure 4 illustrates low correlation between $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$ during unlearning, supporting the reasonabaleness of Assumption 3.2. Note that the results are obtained using SGD optimizer.
C.2 Additional Results of Image Classification
@@
Table 12: Additional performance summary of MU methods for image classification. (a) $50\%$ random subset of CIFAR-10, (b) $10\%$ random subset of CIFAR-100 and (c) $10\%$ random subset of SVHN. ResNet-18 is used. All results are presented as mean and standard deviation across $5$ trials with different random data. Performance gap from RT are indicated in blue.
@@
C.3 Additional Results of Image Generation
@@
C.4 Additional Results of Large Language Models
@@
C.5 Additional Ablation Studies
@@
Ablation studies on Adam-trained ResNet-18 and SCRUB [8]. We conduct ablation studies based on the configurations of Adam-trained ResNet-18 + SFRon and SGD-trained ResNet-18 + SCRUB. As illustrated in Table 17 and 18, Adam (F) + Adam (R) is the best configuration in both cases, indicating that the optimal optimizer for retaining is influenced by the choice of optimizer during pretraining and the specific MU algorithms employed.
@@
Shared optimizer with larger momentum coefficient. We further compare DualOptim with shared optimizers that employ larger momentum coefficients. As observed in Table 19, while increasing the momentum coefficient $\alpha$ can reduce performance variation when using SGD, it also results in slower convergence and ultimately leads to suboptimal performance. In contrast, increasing $\beta_{1}$ has only a marginal impact when using Adam. This is because Adam’s momentum update rule, i.e. ${\bm{m}}=\beta_{1}{\bm{m}}+(1-\beta_{1})\widehat{{\bm{g}}}$, inherently provides a stronger smoothing effect compared to SGD, i.e. ${\bm{m}}=\alpha{\bm{m}}+\widehat{{\bm{g}}}$, when the same momentum coefficient is applied.
@@
C.6 Unlearning Process of Other MU Methods
The unlearning processes of SalUn [6] and SCRUB [8] are illustrated in Figure 6 and 7, respectively. Compared to SFRon (see in Figure 2), SalUn and SCRUB exhibits less performance fluctuation. However, they underperform SFRon by a large margin. Despite that, our method is still effective in boosting their unlearning performance.
(a) Forget Accuracy (FA)
(b) Retain Accuracy (RA)
(c) Test Accuracy (TA)
(d) MIA
(a) Forget Accuracy (FA)
(b) Retain Accuracy (RA)
(c) Test Accuracy (TA)
(d) MIA

### v2

@@ B.1 Implementation Details for Image Classification
For CIFAR-10, CIFAR-100, and SVHN using ResNet-18, all baselines use the SGD optimizer with momentum of $0.9$, weight decay of $5\times 10^{-4}$, and batch size of $128$ if not specified. For TinyImageNet, Swin-T, the models are initialized from torchvision weight pre-trained on ImageNet. All baselines use the Adam optimizer with a weight decay of $5\times 10^{-4}$ and batch size of $128$ if not specified. Summaries of the hyperparameters for each method on each dataset are shown in Table 7 - 10. Note that the unspecified hyperparameters are the same as the default ones reported in their original papers.
@@
Pretrain epoch =200=200, cosine scheduler, $\eta=0.1$
RT epoch =200=200, cosine scheduler, $\eta=0.1$
FT epoch =10=10, constant scheduler, $\eta=0.01$
GA epoch =10=10, constant scheduler, $\eta=5\times 10^{-4}$
RL epoch =10=10, constant scheduler, $\eta_{f}=0.01$, $\eta_{r}=0.01$
SCRUB epoch =10=10, constant scheduler, Adam, $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch =10=10, constant scheduler, $\eta_{f}=0.018$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch =10=10, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=1.5\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch =10=10, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=1.3\times 10^{-4}$, $\eta_{r}=0.01$
@@
Pretrain epoch =200=200, cosine scheduler, $\eta=0.1$
RT epoch =200=200, cosine scheduler, $\eta=0.1$
FT epoch =10=10, constant scheduler, $\eta=0.01$
GA epoch =10=10, constant scheduler, $\eta=8\times 10^{-5}$
RL epoch =10=10, constant scheduler, $\eta_{f}=0.01$, $\eta_{r}=0.01$
SCRUB epoch =10=10, constant scheduler, Adam, $\eta_{f}=1.6\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch =10=10, constant scheduler, $\eta_{f}=2.5\times 10^{-4}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch =10=10, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2.8\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch =10=10, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=5.5\times 10^{-5}$, $\eta_{r}=0.01$
@@
Pretrain epoch =200=200, cosine scheduler, $\eta=0.1$
RT epoch =200=200, cosine scheduler, $\eta=0.1$
FT epoch =10=10, constant scheduler, $\eta=4.5\times 10^{-2}$
GA epoch =10=10, constant scheduler, $\eta=5.2\times 10^{-4}$
RL epoch =10=10, constant scheduler, $\eta_{f}=8\times 10^{-4}$, $\eta_{r}=0.01$
SCRUB epoch =10=10, constant scheduler, Adam, $\eta_{f}=6\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch =10=10, constant scheduler, $\eta_{f}=1.3\times 10^{-3}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch =10=10, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch =10=10, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=1.9\times 10^{-4}$, $\eta_{r}=0.01$
@@
Pretrain epoch =200=200, cosine scheduler, $\eta=0.1$
RT epoch =200=200, cosine scheduler, $\eta=0.1$
FT epoch =10=10, constant scheduler, $\eta=2.2\times 10^{-2}$
GA epoch =10=10, constant scheduler, $\eta=5.2\times 10^{-4}$
RL epoch =10=10, constant scheduler, $\eta_{f}=8\times 10^{-3}$, $\eta_{r}=0.01$
SCRUB epoch =10=10, constant scheduler, Adam, $\eta_{f}=8\times 10^{-6}$, $\eta_{r}=1\times 10^{-4}$
SalUn epoch =10=10, constant scheduler, $\eta_{f}=1.25\times 10^{-2}$, $\eta_{r}=0.01$
@@
SCRUB+DO epoch =10=10, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2.2\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SalUn+DO epoch =10=10, constant scheduler, Adam (F) + SGD (R), $\eta_{f}=2.5\times 10^{-6}$, $\eta_{r}=0.01$
@@
Pretrain epoch =10=10, cosine scheduler, $\eta=1\times 10^{-4}$
RT epoch =10=10, cosine scheduler, $\eta=1\times 10^{-4}$
FT epoch =1=1, constant scheduler, $\eta=1\times 10^{-4}$
GA epoch =1=1, constant scheduler, $\eta=5\times 10^{-6}$
RL epoch =1=1, constant scheduler, $\eta_{f}=5\times 10^{-5}$, $\eta_{r}=1\times 10^{-4}$
SCRUB epoch =1=1, constant scheduler, Adam, $\eta_{f}=2\times 10^{-5}$, $\eta_{r}=2\times 10^{-5}$
SalUn epoch =1=1, constant scheduler, $\eta_{f}=9\times 10^{-5}$, $\eta_{r}=9\times 10^{-5}$
@@
SCRUB+DO epoch =1=1, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=2\times 10^{-5}$, $\eta_{r}=2\times 10^{-5}$
SalUn+DO epoch =1=1, constant scheduler, Adam (F) + Adam (R), $\eta_{f}=9\times 10^{-5}$, $\eta_{r}=9\times 10^{-5}$
@@ B.3 Implementation Details for Natural Language Processing
Phi-1.5 ME+GD+DO epoch =5=5, $\eta_{f}=1\times 10^{-5}$, $\eta_{r}=1\times 10^{-5}$
IDK+AP+DO epoch =5=5, $\eta_{f}=8\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
LLaMA 2 ME+GD+DO epoch =5=5, $\eta_{f}=1.45\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
IDK+AP+DO epoch =5=5, $\eta_{f}=1.25\times 10^{-6}$, $\eta_{r}=1\times 10^{-5}$
LLaMA 2 LoRA ME+GD+DO epoch =5=5, $\eta_{f}=1\times 10^{-4}$, $\eta_{r}=2\times 10^{-4}$
IDK+AP+DO epoch =5=5, $\eta_{f}=5\times 10^{-5}$, $\eta_{r}=2\times 10^{-4}$
@@ Appendix C Additional Results
C.1 Non-positive Correlation between Forget and Retain Gradients
@@
Figure 4 illustrates non-positive correlation between $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$ during unlearning, supporting the reasonableness of Assumption 3.2. Note that the results are obtained using SGD optimizer.
C.2 Unlearning Process of Other MU Methods
The unlearning processes of SalUn [6] and SCRUB [8] are illustrated in Figure 5 and 6, respectively. Compared to SFRon (see in Figure 2), SalUn and SCRUB exhibits less performance fluctuation. However, they underperform SFRon by a large margin. Despite that, our method is still effective in boosting their unlearning performance.
(a) Forget Accuracy (FA)
(b) Retain Accuracy (RA)
(c) Test Accuracy (TA)
(d) MIA
(a) Forget Accuracy (FA)
(b) Retain Accuracy (RA)
(c) Test Accuracy (TA)
(d) MIA
C.3 Additional Results of Image Classification
@@
Table 13: Additional performance summary of MU methods for image classification. (a) $50\%$ random subset of CIFAR-10, (b) $10\%$ random subset of CIFAR-100 and (c) $10\%$ random subset of SVHN. ResNet-18 is used. All results are presented as mean and standard deviation across 55 trials with different random data. Performance gap from RT are indicated in blue.
@@
C.4 Additional Results of Image Generation
@@
C.5 Additional Results of Large Language Models
@@
(a) untargeted unlearning
(b) targeted unlearning
C.6 Additional Ablation Studies
@@
Ablation studies on Adam-trained ResNet-18 and SCRUB [8]. We conduct ablation studies based on the configurations of Adam-trained ResNet-18 + SFRon and SGD-trained ResNet-18 + SCRUB. As illustrated in Table 18 and 19, Adam (F) + Adam (R) is the best configuration in both cases, indicating that the optimal optimizer for retaining is influenced by the choice of optimizer during pretraining and the specific MU algorithms employed.
@@
Shared optimizer with larger momentum coefficient. We further compare DualOptim with shared optimizers that employ larger momentum coefficients. As observed in Table 20, while increasing the momentum coefficient $\alpha$ can reduce performance variation when using SGD, it also results in slower convergence and ultimately leads to suboptimal performance. In contrast, increasing $\beta_{1}$ has only a marginal impact when using Adam. This is because Adam’s momentum update rule, i.e. $^{\bm{m}}=\beta_{1}{\bm{m}}+(1-\beta_{1})\widehat{{\bm{g}}}$, inherently provides a stronger smoothing effect compared to SGD, i.e. $^{\bm{m}}=\alpha{\bm{m}}+\widehat{{\bm{g}}}$, when the same momentum coefficient is applied.